OpenStudy (anonymous):

Determine whether a triangle can be formed with the given side lengths. If so, use Heron's formula to find the area of the triangle. a = 240 b = 132 c = 330

OpenStudy (anonymous):

see if sum of any two sides > third side.

OpenStudy (anonymous):

2s=a+b+c \[area of \triangle=\sqrt{s \left( s-a \right)\left( s-b \right)\left( s-c \right)}\] now you can solve.